OpenStudy (anonymous):

A and B are independent events. Find P(A and B) 1. P(A)=1/6 , P(B) = 2/5 2. P(A)=9/20 , P(B) =3/4

jimthompson5910 (jim_thompson5910):

If A and B are independent events, then P(A and B) = P(A) * P(B)

jimthompson5910 (jim_thompson5910):

For #1, P(A and B) = P(A) * P(B) P(A and B) = (1/6)*(2/5) P(A and B) = (1*2)/(6*5) P(A and B) = 2/30 P(A and B) = 1/15 I'll let you do #2